# ComGPT: Detecting Local Community Structure with Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs), like GPT, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks, such as node classification. Despite their strong abilities in graph reasoning tasks, they lack specific domain knowledge and have a weaker understanding of community-related graph information, which hinders their capabilities in the community detection task. Moreover, local community detection algorithms based on seed expansion, referred to as seed expansion algorithms, often face the seed-dependent problem, community diffusion, and free rider effect. To use LLMs to overcome the above shortcomings, we explore a GPT-guided seed expansion algorithm named ComGPT. ComGPT iteratively selects potential nodes by local modularity $M$ from the detected community’s neighbors, and subsequently employs LLMs to choose the node to join the detected community from these selected potential nodes. To address the above issues faced by LLMs, we improve graph encoding method, called Incident, by incorporating community knowledge to improve LLMs’s understanding of community-related graph information. Additionally, we design the NSG (Node Selection Guide) prompt to enhance LLMs’ understanding of community characteristics. Experimental results demonstrate that ComGPT outperforms the comparison methods, thereby confirming the effectiveness of the improved graph encoding method and prompts.
Introduction
@@
Identifying community structures from networks has wide applications, such as identifying protein complexes (Yin et al. 2017). Local community detection has gained attention due to its ability to quickly identify the community containing a specific node (Ni et al. 2024a, b). Various methods are designed for detecting local community, including seed expansion algorithms (Luo, Wang, and Promislow 2008), heat diffusion techniques (Kloster and Gleich 2014), and non-negative matrix factorization algorithms (Wang et al. 2011). Among these methods, the seed expansion algorithms have garnered widespread acclaim for their versatility and ease of implementation. The seed expansion algorithms typically enlarge the community by selecting a node with the high scoring function value at each step (Moradi, Olovsson, and Tsigas 2014).
@@
However, seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013), shown in Figure 1. The seed-dependent problem refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by nodes 5 and 1 in Figure 1(a). The community diffusion problem means that the detected community contains nodes from different communities, as illustrated by node 1 in Figure 1(b). The free rider effect refers to mistakenly adding unrelated communities to enhance the scoring function value of a detected community, as illustrated by node 7 in Figure 1(c). To address these problems, researchers developed different strategies such as identifying core node sets (Guo et al. 2022) and designing scoring functions with a bias towards nodes with large degrees (Luo et al. 2018).
((a))
((b))
((c))
Figure 1: Some shortcomings faced by seed expansion algorithms. Nodes in the same community are marked with the same color.
In this study, inspired by the success of Large Language Models (LLMs) in many fields (Gu et al. 2024; Zong and Krishnamachari 2024; Yang et al. 2022)(Adhikari et al. 2020; Ammanabrolu and Riedl 2021)(Wang et al. 2023), we explore an alternative approach based on LLMs, called ComGPT (GPT-guided seed expansion algorithm). ComGPT leverages LLMs to guide the local community detection process to alleviate the seed-dependent problem, community diffusion, and free rider effect. The one-question-one-answer modality of LLMs aligns well with the incremental nature of seed expansion algorithms. It bridges the current gap in LLMs applicability to the community detection domain.
Specifically, we leverage LLMs to select nodes to join the community instead of the scoring function commonly used in seed expansion algorithms. However, the application of LLMs to select nodes faces the following challenges: 1) Lack of graph encoding methods for providing LLMs with sufficient community-related graph information. To address this, we improve the graph encoding method, called Incident (Fatemi, Halcrow, and Perozzi 2023), by incorporating community information. 2) Lack of domain knowledge for LLMs in dealing with local community detection. It limits the capabilities of LLMs. To address this, we design prompts to assist LLMs in understanding community definitions and how to select nodes. Additionally, LLMs receive only a partial network as input, not the entire network, to minimize costs by reducing the number of tokens processed.
@@
  • We design an improved version of the typical graph encoding method, named Incident (Fatemi, Halcrow, and Perozzi 2023), and NSG (Node Selection Guide) prompt. This improved version incorporates both graph topology and community information, which enhances LLMs to understand the current state of the detected community. NSG prompt that integrates community features facilitates LLMs’s comprehension of domain knowledge.
  • The results on real datasets indicate that ComGPT outperforms comparison methods. The ablation study experiments confirm that LLMs play a crucial role in the ComGPT’s performance. Furthermore, prompt comparison experiments reveal that the designed prompt effectively aids LLMs in understanding the graph structure and community situation.
Related work
Local community detection
Local community detection focuses on discovering the community to which the seed node belongs (Ni et al. 2024c, 2023). A variety of methods, like $k$-core decomposition (Cui et al. 2014) and seed expansion algorithms (Clauset 2005), are developed for detecting local communities. Seed expansion algorithms expand communities by selecting nodes with high scoring function values (Moradi, Olovsson, and Tsigas 2014). For example, the R method (Clauset 2005) and M method (Luo, Wang, and Promislow 2008) choose nodes based on the local modularity R and M, respectively. The seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013). To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets (Guo et al. 2022; Ji, Guo, and Yu 2022), finding the nearest nodes with greater centrality (Luo et al. 2020), alternating strategy of strong fusion and weak fusion (Shang et al. 2023), and designing membership functions with a bias towards nodes with large degrees (Luo et al. 2018). Our work mitigates the aforementioned problems by utilizing LLMs, distinguishing them from existing studies.
LLMs in graph reasoning
In recent years, researchers have explored the application of LLMs in graph reasoning (Zhang 2023; Chen et al. 2023). Wang et al. demonstrated LLMs’ graph reasoning abilities and designed the NLGraph dataset (Wang et al. 2023) to evaluate these capabilities. Furthermore, Guo et al. emphasized the design and improvement of graph encoding techniques, which critically affect LLMs’ interpretation of input data and its subsequent output (Guo et al. 2023). Therefore, Adjacency and Incident (Fatemi, Halcrow, and Perozzi 2023) are designed to encode graph structured data into text for LLMs. Adjacency uses integer node encoding and parenthesis edge encoding. Incident uses integer node encoding and incident edge encoding. Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection. Our work introduces a novel graph encoding method based on Incident (Fatemi, Halcrow, and Perozzi 2023), tailored for the community detection domain, that incorporates community structure and distinguishes itself from existing approaches.
Furthermore, various prompts are designed for specific tasks to make LLMs better adapt to these tasks (Sahoo et al. 2024), such as the general CoT (Chain-of-Thought) prompt (Wei et al. 2022) and the BaG (Build a Graph) prompt (Wang et al. 2023) for graph reasoning tasks. However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks. This paper designs a prompt, named NSG, suitable for the field of local community detection, which is different from the above prompts.
Approach
@@
We introduce a GPT-guided seed expansion algorithm named ComGPT, based on GPT-3.5-turbo. This section begins with an overview of ComGPT, followed by detailed descriptions of each step of ComGPT.
Overview
@@
3:     $PN\leftarrow$ potential nodes identification with $C$
@@
5:        $gtxt\leftarrow$ graph encoding with ($C\cup PN$, $C$, $PN$)
6:        $v\leftarrow$ node selection by GPT-3.5-turbo with ($gtxt$, NSG, Qsel )
@@
12:     $gtxt\leftarrow$ graph encoding with ($C\cup N$, $C$, $PN$)
13:     $v\leftarrow$ node supplementation by GPT-3.5-turbo with ($gtxt$, NSU, Qsup)
@@
As shown in Figure 2, ComGPT contains four main steps. To better understand the interactions among these steps, Algorithm 1 shows the overall process of ComGPT. Initially, community $C$ contains seed node $v_{s}$ and $cands$ is empty (line 1), where $cands$ stores the community whenever the potential node identification step returns an empty set. First, ComGPT selects potential nodes $PN$ with the step of potential node identification (line 3), used for subsequent graph encoding. In order for GPT-3.5-turbo to understand the graph structure and community information, the local network, i.e., the subgraph consisting of the nodes in both current community $C$ and $PN$ as well as edges between these nodes, is converted into text, i.e., $gtxt$ (line 5). Utilizing $gtxt$, NSG (Node Select Guide) prompt, and Qsel, GPT-3.5-turbo selects node $v$ from potential nodes $PN$ and adds it to the current community $C$ (lines 6-7). The above steps are repeated until $PN$ is empty. If $PN$ is empty, the community $C$ is added to the $cands$ (line 9). The fact that $PN$ is empty does not necessarily imply that the neighboring nodes of the community are weakly connected to the community. Therefore, based on $gtxt$, NSU (Node SUpplementatio) prompt, and Qsup, ComGPT continues to identify nodes through node supplementation (lines 10-13). If GPT-3.5-turbo identifies suitable node $v$ to add to the community, then ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 15). Otherwise, GPT-3.5-turbo returns “null” and ComGPT stops community expansion (line 17). After the above process is completed, ComGPT selects the community with the highest local modularity $M$ calculated by formula (1) from $cands$ as the final community (line 20).
Potential nodes identification
To mitigate the cost associated with processing nodes through GPT-3.5-turbo, ComGPT identifies potential nodes that may belong to the current community and provides these nodes to GPT-3.5-turbo for selection. Different scoring functions are designed in the seed expansion algorithms to select nodes (Moradi, Olovsson, and Tsigas 2014), which can be used to choose potential nodes, such as modularity $M$ (Luo, Wang, and Promislow 2008), $R$ (Clauset 2005), and $Q_{lcd}$ (Shang et al. 2023). ComGPT uses local modularity $M_{C}$ (Luo, Wang, and Promislow 2008) to select potential nodes. $M_{C}$ is calculated as,
@@
Figure 3: An toy example of graph encoding. The middle part illustrates the text corresponding to graph topological structure. The lower part illustrates the text corresponding to supplementary community information.
@@
The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT-3.5-turbo. The graph data includes two aspects. 1) Graph topological structure. We adopt Incident (Fatemi, Halcrow, and Perozzi 2023) to let GPT-3.5-turbo understand the graph’s topological structure. Incident represents the graph topology by introducing the first-order neighbors of each node. To reduce computational overhead, the graph structure input to GPT-3.5-turbo is a local network. For the node selection by GPT-3.5-turbo step, the local network comprises nodes in $C\cup PN$ and the edges between them. For the node supplementation by GPT-3.5-turbo step, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$. 2) Supplementary knowledge. We describe node information from a community perspective to enhance GPT-3.5-turbo’s understanding of community structure. This step involves outlining which nodes are within the community, those that are outside, and the connections of outside nodes to both within and outside the community. The outside nodes are potential nodes in $PN$ from the step of potential nodes identification. The text corresponding to the topological structure and the text corresponding to the supplementary information are concatenated to form $gtxt$. To better understanding, an example is given in Figure 3.
Node selection by GPT-3.5-turbo
@@
Node selection refers to choosing the most suitable node for community expansion, based on NSG prompt, graph text $gtxt$, and question Qsel. Graph text $gtxt$ is obtained by the step of graph encoding with the local network, $C$, and $PN$, where the local network consists of the nodes in $C\cup PN$ and edges between these nodes. Below, we introduce NSG prompt and question Qsel.
@@
  • Prompt 1: The more an outside node is connected to other outside nodes, the higher the likelihood of its selection. The outside nodes provided to GPT-3.5-turbo are potential nodes from the step of potential nodes identification. This prompt could alleviate the seed-dependent problem and the free rider effect.
@@
    The reason for the seed-dependent problem is low-degree node preference. Take seed node 5 in Figure 1(a) and M method (Luo, Wang, and Promislow 2008) as an example to illustrate it. Nodes 1, 7, and 11 are each connected to the node in the community [5] by an edge, and node 11, having the smallest degree, is selected so that the community has the fewest external edges. Thus, the M method (Luo, Wang, and Promislow 2008) selects node 11 to join the community. Nodes that connect to many other potential nodes are usually high-degree nodes, which help to avoid bias towards low-degree nodes and thereby alleviate the seed-dependent problem. Nodes 1, 7, and 11 are potential nodes for community [5]. According to prompt 1, node 1 will be selected over the mistakenly considered node 11.
    The reason for the free rider effect is adding nodes that are sparsely connected to the community to the community. Nodes that are connected to multiple potential nodes indirectly maintain close connections to the community, which alleviates the free rider effect. Taking Figure 1(c) as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes. According to prompt 1, nodes 6 and 8 have a higher probability of being selected. Node 4 is a node that may cause the free rider effect, and will not be selected.
@@
  • Prompt 2: Prioritize selecting outside nodes that are connected to multiple nodes within the community. The outside nodes mentioned in prompts are potential nodes from the step of potential nodes identification. This prompt could alleviate the community diffusion problem. The reason for the community diffusion is the insufficient consideration of the connections between nodes to be added to the community and the nodes within the community. Therefore, the prompt considers the connection between potential nodes and the community, thereby alleviating the community diffusion problem. Taking Figure 1(b) as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where nodes 13 or 18 may lead to community diffusion. According to prompt 2, nodes 2 and 5 have a higher probability of being selected, rather than mistakenly selecting nodes 13 or 18.
    The first prompt is to consider the closeness of the node to potential nodes outside the community, and the second prompt is to consider the closeness of the node to nodes within the community. This is consistent with the characteristics of the community being close inside, thus helping GPT-3.5-turbo understand the characteristics of the nodes to be selected. To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for selecting nodes is designed: “You’re doing local community detection. Based on the graph data and prompt, please select a node that you think is most likely to belong to the current community $C$ for community expansion. Provide a detailed explanation.”. Based on the above prompts, the text $gtxt$ obtained by graph encoding method, and the question Qsel, GPT selects the node $v$ from potential nodes.
Node supplementation by GPT-3.5-turbo
Node supplementation involves evaluating whether to continue the expansion process or terminate it. If continuing, it also determines which additional nodes should be selected. Next, we introduce the information needed for graph encoding, NSU prompt, and question Qsup in turn. For graph encoding, to provide GPT-3.5-turbo with sufficient information to make a decision, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$. Besides, ComGPT forcibly adds the $k$ nodes with the highest $M$ values from the first neighbors of community $C$ to $PN$. NSU prompt is based on the definition of the community and whether these nodes can maintain the cohesion of the community. To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for supplementing nodes is designed: “Please analyze whether these nodes should be added to the community $C$. The probability of not adding nodes is higher. But it doesn’t mean you always refuse to add nodes. If you think there is a suitable node, please output its node number.”. Based on the above prompts, the text obtained by graph encoding method, and the question Qsel, if GPT-3.5-turbo determines that any of nodes warrant inclusion in the community, it returns the node’s number; otherwise, it returns “null”.
Experiments
Experimental settings
 Dataset M R DMF_M DMF_R LCDNN LS ComGPT
Dolphins Jaccard 0.3370 0.3128 0.6270 0.6320 0.6438* 0.7735 0.7035
Fscore 0.4673 0.4478 0.7254 0.7241 0.7558* 0.8385 0.7684
Polbooks Jaccard 0.4266 0.3986 0.6449 0.6437 0.6673 0.6919 0.6468*
Fscore 0.5184 0.4960 0.7294 0.7285 0.7463 0.7764 0.7320*
Football Jaccard 0.6117 0.6134 0.8546 0.8579 0.7547 0.2406 0.7571*
Fscore 0.6706 0.6699 0.8881 0.8893 0.8255* 0.3593 0.8103
Amazon Jaccard 0.6961 0.6771 0.7633 0.7686 0.7665* 0.6506 0.7734
Fscore 0.7706 0.7554 0.8353 0.8361* 0.8372 0.7420 0.8421
DBLP Jaccard 0.3239 0.3207* 0.2945 0.2928 0.2208 0.0650 0.3259
Fscore 0.4185 0.4185 0.4032 0.4009 0.3152 0.0850 0.4359
Table 1: Experimental results on five datasets. On each dataset, the highest Fscore (Jaccard) is highlighted in bold, the second highest one is highlighted by underline, and the third highest one is highlighted with a superscript asterisk.
Datasets

### v3

@@ Abstract
Large Language Models (LLMs), like GPT-3.5-turbo, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks, such as node classification. Despite their strong abilities in graph reasoning tasks, they lack specific domain knowledge and have a weaker understanding of community-related graph information, which hinders their capabilities in the community detection task. Moreover, local community detection algorithms based on seed expansion, referred to as seed expansion algorithms, often face several shortcomings, including the seed-dependent problem, community diffusion, and free rider effect. To use LLMs to overcome the above shortcomings, we explore a GPT-guided seed expansion algorithm named ComGPT. ComGPT iteratively selects potential nodes by local modularity from the detected community’s neighbors, and subsequently employs LLMs to choose the node from these selected potential nodes to join the detected community. To improve LLMs’ understanding of community-related graph information, we propose ComIncident, a graph encoding method that incorporates community knowledge and is designed for the community detection task. Additionally, we design the Node Selection Guide (NSG) prompt to enhance LLMs’ understanding of community characteristics. Experimental results demonstrate that ComGPT outperforms the baselines, thereby confirming the effectiveness of the ComIncident and the NSG prompt.
I Introduction
@@
Identifying community structures from networks has wide applications, such as identifying protein complexes [1]. Local community detection has gained attention due to its ability to quickly identify the community containing a specific node [2, 3]. Various methods are designed for detecting local community, including seed expansion algorithms [4], pagerank based techniques [5], flow-propagation based techniques [6, 7], and non-negative matrix factorization algorithms [8]. Among these methods, the seed expansion algorithms have garnered widespread acclaim for their versatility and ease of implementation. It typically enlarges the community by selecting a node with the high scoring function value at each step [9].
Figure 1: Seed-dependent problem. The community detected by M method [4] with starting nodes 5 is {5, 11, 6, 7, 17}. This community misses many nodes. Nodes in the same community are marked with the same color.
Figure 2: Community diffusion. The community detected by the M method [4] with node 1 as the starting seed node is {1, 2, 3, 4, 7, 8, 9, 13, 14, 15, 16, 17, 18, 20}, which incorrectly includes nodes from two different communities. Nodes in the same community are marked with the same color.
Figure 3: Free rider effect. The target community consists of nodes 5, 6, 7, 8, and 9. The detected community not only contains nodes in the target community {5, 6, 7, 8, 9} but also erroneously includes all nodes in the unrelated community {1, 2, 3, 4, 10, 11}. Nodes in the same community are marked with the same color.
@@
However, seed expansion algorithms often face the seed-dependent problem [10], community diffusion, and free rider effect [11, 12, 13], which are shown in Figure 1, 2, and 3, respectively. The seed-dependent problem refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by node 5 in Figure 1. The community diffusion problem means that the detected community contains nodes from different communities, as illustrated by node 1 in Figure 2. The free rider effect refers to mistakenly adding unrelated communities to enhance the scoring function value of a detected community, as illustrated by node 7 in Figure 3. To address these problems, researchers developed different strategies such as identifying core node sets [14], finding the nearest nodes with greater centrality [15], alternating strategy of strong fusion and weak fusion [16], and so on.
The above studies rely on predefined heuristic rules, which may fail to adapt to diverse structures across networks. Recently, Large Language Models (LLMs) have emerged as powerful reasoning engines [17, 18, 19, 20, 21, 22]. Unlike these heuristic-based approaches, LLMs do not rely on a fixed set of rules and can be prompted to reason about diverse network structures. Therefore, we explore a GPT-guided seed expansion approach based on LLMs, called ComGPT. ComGPT leverages LLMs to guide the detection of a local community. The one-question-one-answer modality of LLMs aligns well with the incremental nature of seed expansion algorithms, making it naturally suitable for this task. This approach bridges the current gap in LLMs applicability to the community detection domain.
Specifically, we leverage LLMs to select nodes to join the community instead of the scoring function used in seed expansion algorithms. However, the application of LLMs to select nodes faces the following challenges:
  1. 1.
    Lack of graph encoding methods for providing LLMs with sufficient community-related graph information. To address this, we propose ComIncident (Community-Enriched Incident), a variant of the graph encoding method Incident [23], which incorporates community information.
  2. 2.
    Lack of domain knowledge for LLMs in dealing with local community detection. It limits the capabilities of LLMs. To address this, we design prompts to assist LLMs in understanding community definitions and how to select nodes.
  3. 3.
    Token count constraints. To address this, LLMs receive only a partial network as input, not the entire network, to minimize costs by reducing the number of tokens processed.
@@
  2. 2.
    We design an graph encoding method, named ComIncident, and NSG (Node Selection Guide) prompt. ComIncident is the first graph encoding method designed for the community detection task, which enhances LLMs to understand the current state of the detected community. The NSG prompt, which integrates community features, facilitates LLMs’s comprehension of domain knowledge.
  3. 3.
    The results on real datasets indicate that ComGPT outperforms baselines. The ablation study experiments confirm that LLMs play a crucial role in the ComGPT’s performance. Furthermore, prompt comparison experiments reveal that the designed prompt effectively aids LLMs in understanding the graph structure and community situation.
The rest of the paper is organized as follows. Section II presents the proposed algorithm. Section III reports experimental results. Section IV summarizes the related works. Section V concludes the paper.
II Approach
@@
We introduce a GPT-guided seed expansion algorithm named ComGPT. This section begins with an overview of ComGPT, followed by detailed descriptions of each step of ComGPT. Table I lists the important notations used in the paper.
II-A Overview
TABLE I: Important notations
Notation Definition
$G$ network
$C$ community
NN neighbor node of $C$
$v_{s}$ seed node
PNPN potential node
candscands communities each time $PN$ is empty
gtxtgtxt graph text obtained by the graph encoding step
NSGNSG node selection guide prompt
NSUNSU node supplementation prompt
$Q_{sel}$ node selection question
$Q_{sup}$ node supplementation question
As shown in Figure 4, ComGPT contains four main steps: potential node identification, graph encoding, node selection by GPT-3.5-turbo, and node supplementation by GPT-3.5-turbo. The idea of ComGPT uses potential node identification step (Section II-B) to select potential nodes from the neighbors of the current community. Then, it uses node selection by GPT-3.5-turbo step (Section II-D) and node supplementation by GPT-3.5-turbo step (Section II-E) selects nodes to join the community. Since GPT-3.5-turbo cannot directly process graph structures and community structures, the graph encoding step (Section II-C) is used to encode the graph structure and community structure into text that GPT-3.5-turbo can understand.
Algorithm 1 shows the overall process of ComGPT. Initially, community $C$ contains seed node $v_{s}$ and candscands is empty (line 1), where candscands stores the communities each time $PN$ is empty. First, ComGPT selects potential nodes $PN$ with the step of potential node identification (line 3), used for subsequent graph encoding. In order for GPT-3.5-turbo to understand the graph structure and community information, the local network, i.e., the subgraph consisting of the nodes in both current community $C$ and $PN$ as well as edges between these nodes, is converted into text, i.e., gtxtgtxt (line 5). Utilizing gtxtgtxt, GPT-3.5-turbo selects node $v$ from potential nodes $PN$ and adds it to the current community $C$ (lines 6-7). The above steps are repeated until $PN$ is empty. Then the community $C$ is added to the candscands (line 9). $PN$ being empty does not necessarily imply that the neighboring nodes of the community are weakly connected to it. Therefore, ComGPT continues to identify nodes through node supplementation (lines 10-13). If GPT-3.5-turbo identifies suitable node $v$, ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 15). Otherwise, GPT-3.5-turbo returns “null” and ComGPT stops community expansion (line 17). After the above process is completed, ComGPT selects the community with the highest local modularity $M$ calculated by formula (1) from candscands as the final community (line 20). During the expansion process, if no new node is selected in potential nodes identification step (Section II-B), i.e., $PN$ is empty (line 3), ComGPT jumps out of the inner while loop. Thereafter, if the node selected in the node supplementation by GPT-3.5-turbo is null (line 14), ComGPT terminates.
@@ II-A Overview
3:  $PN\leftarrow$ potential nodes identification with $C$ // Section II-B
@@
5:   $gtxt\leftarrow$ graph encoding with ($\cup PN, CC$, PNPN) // Section II-C
6:   $v\leftarrow$ node selection by GPT-3.5-turbo with (gtxtgtxt, NSG, Qsel ) // Section II-D
@@
12:  $gtxt\leftarrow$ graph encoding with ($\cup N, CC$, PNPN) // Section II-C
13:  $v\leftarrow$ node supplementation by GPT-3.5-turbo with (gtxtgtxt, NSU, Qsup) // Section II-E
@@
II-B Potential nodes identification
To mitigate the cost associated with processing nodes through GPT-3.5-turbo, ComGPT identifies potential nodes $PN$ that may belong to the current community and provides $PN$ to GPT-3.5-turbo for selection. Different scoring functions are designed in the seed expansion algorithms to select nodes [9], which can be used to choose potential nodes, such as modularity $M$ [4], $R$ [24], and $Q_{lcd}$ [16]. ComGPT uses local modularity $M_{C}$ [4] to select potential nodes. $M_{C}$ is calculated as,
@@ II-C Graph encoding
Figure 5: An toy example of graph encoding. The middle part illustrates the text corresponding to topology graph. The lower part illustrates the text corresponding to supplementary knowledge.
@@
The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT-3.5-turbo. We design graph encoding method, called ComIncident, for the community detection task. ComIncident includes two aspects: 1) Topology graph. The graph topology is represented by introducing the first-order neighbors of each node, as in Incident [23]. 2) Supplementary knowledge. We describe node information from a community perspective to enhance GPT-3.5-turbo’s understanding of community structure. It involves outlining which nodes are within the community, those that are outside, and the connections of potential nodes to both within and outside the community. The topological structure of local network and the supplementary information are merged to form gtxtgtxt. ComIncident is an improved version of Incident. Incident utilizes only the topology graph, whereas ComIncident incorporates both the topology graph and supplementary knowledge. To better understand ComIncident, an example is given in Figure 5.
For the node selection by GPT-3.5-turbo step (Section II-D), the local network comprises nodes in $C\cup PN$ and the edges between them. For the node supplementation by GPT-3.5-turbo step (Section II-E), the local network consists of nodes in $C\cup N$ and edges between these nodes, where $N$ includes the first and second-order neighbors of community $C$.
II-D Node selection by GPT-3.5-turbo
@@
Node selection refers to choosing the most suitable node for community expansion, based on NSG (Node Selection Guide) prompt, graph text gtxtgtxt, and question Qsel. Graph text gtxtgtxt is obtained by the graph encoding step (Section II-C) with the local network, $C$, and $PN$, where the local network consists of the nodes in $C\cup PN$ and edges between these nodes. Below, we introduce the NSG prompt and question Qsel, where the NSG prompt contains NSG_1 and NSG_2.
@@
Prompt NSG_1: The more an outside node is connected to other outside nodes, the higher the likelihood of its selection.
@@
The outside nodes provided to GPT-3.5-turbo are potential nodes from the potential nodes identification step (Section II-B). This prompt could alleviate the seed-dependent problem and the free rider effect. The reason for the seed-dependent problem is low-degree node preference. Take seed node 5 in Figure 1 and M method [4] as an example to illustrate it. Nodes 1, 7, and 11 are connected to nodes within the community [5]. The M method [4] selects node 11, having the smallest degree, to join the community, ensuring the fewest external edges of the community. With the incorporation of the prompt NSG_1, nodes that connect to many other potential nodes are prioritized. Consequently, node 1 is selected over the mistakenly considered node 11. That is, prompt NSG_1 helps to avoid bias towards low-degree nodes and thereby alleviates the seed-dependent problem. The reason for the free rider effect is adding nodes that are sparsely connected to the community to the community. Nodes that are connected to multiple potential nodes indirectly maintain close connections to the community, which alleviates the free rider effect. Taking Figure 3 as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes. According to NSG_1, nodes 6 and 8 have a higher probability of being selected. Node 4 is a node that may cause the free rider effect, and will not be selected.
@@
Prompt NSG_2: Prioritize selecting outside nodes that are connected to multiple nodes within the community. The outside nodes mentioned in prompts are potential nodes from the potential nodes identification step (Section II-B). This prompt could alleviate the community diffusion. The reason for the community diffusion is the insufficient consideration of the connections between nodes to be added to the community and the nodes within the community. Therefore, the prompt considers the connection between potential nodes and the community, thereby alleviating the community diffusion problem. Taking Figure 2 as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where nodes 13 or 18 may lead to community diffusion. According to NSG_2, nodes 2 and 5 are more likely selected, rather than mistakenly selecting nodes 13 or 18.
The NSG_1 considers the closeness of the node and potential nodes that are highly likely to be added into the community as ComGPT proceeds. The NSG_2 focuses on the closeness of the node and nodes within the community. These two prompts align with the characteristics of communities where nodes in the community are tightly connected, thus helping GPT-3.5-turbo understand the characteristics of the nodes to be selected. To guide GPT-3.5-turbo to provide relevant answers, the question Qsel is designed:
You’re doing local community detection. Based on the graph data and prompt, please select a node that you think is most likely to belong to the current community $C$ for community expansion. Provide a detailed explanation.”.
Based on the above prompts, the graph text gtxtgtxt (Section II-C), and the question Qsel, GPT-3.5-turbo selects the node $v$ from potential nodes.
II-E Node supplementation by GPT-3.5-turbo
Node supplementation involves evaluating whether to continue the expansion process or terminate it. If continuing, it also determines which additional nodes should be selected. Next, we introduce NSU (Node SUpplementation) prompt and question Qsup in turn.
NSU prompt is based on the definition of the community and whether these nodes can maintain the cohesion of the community, as follows:
Definition of community: nodes in the same community are tightly connected, while the nodes in different communities are sparsely connected. The more a node connects with nodes within a community, the more it can increase the community’s cohesion. Conversely, the more it connects with nodes outside the community, the more it can decrease the community’s cohesion. If you choose to add a node to the community, it should make the connections within the community tighter.”.
The question Qsel for supplementing nodes is designed as:
Please analyze whether these nodes should be added to the community $C$. The probability of not adding nodes is higher. But it doesn’t mean you always refuse to add nodes. If you think there is a suitable node, please output its node number.”.
Based on the NSU prompt, the text obtained by the graph encoding step (Section II-C), and the question Qsel, GPT-3.5-turbo determines whether any of the nodes warrant inclusion in the community. If so, it returns the node’s number; otherwise, it returns ‘null’.
ComGPT aims to explore the application of LLMs to local community detection. Specifically, it skillfully combines LLMs with the seed expansion algorithm. Different from traditional seed expansion methods that rely on predefined and rigid heuristic rules, ComGPT leverages the implicit semantic reasoning capability of LLMs to guide the seed expansion process. Notably, we do not simply replace a certain rule of the seed expansion algorithm with LLMs. Instead, ComGPT provides the LLM with guiding principles rather than explicit, step-by-step rules. This leverages the LLM’s inherent reasoning capability to grasp the nuanced semantics of concepts like ‘tightness’ and ‘community membership’. Consequently, the LLMs can be prompted to reason about diverse network structures.
II-F Complexity analysis
The time complexity analysis is as follows. Assuming there are $n$ nodes in the local community and the average degree of nodes is $d$. The potential node identification step involves traversing the neighbors of nodes within the local community and calculating modularity M, with a time complexity of O($nd^{2}$). The graph encoding step requires generating textual descriptions for the current community and potential nodes, with a time complexity proportional to the number of nodes and their degrees, i.e., O($nd$). The core of the node selection and node supplementation steps involves calling the GPT-3.5-turbo API. Since there is no literature defining the time complexity of GPT-3.5-turbo API calls, this portion of the time complexity remains unestimated. ComGPT repeats the above steps $n$ times to get the final community. Excluding the GPT-3.5-turbo API calls, the time complexity of ComGPT is O($n^{2}d^{2}$).
The space complexity analysis is as follows. The potential node identification step stores local community, neighbors of local community, and $k$ potential nodes. In ComGPT, $k$ is treated as a constant with a default value of 10; therefore, the space complexity of this step is O($n+nd+k$). The graph encoding step requires storing the current community along with its potential nodes and the connections between nodes, resulting in a space complexity of O($nd+kd$). The node selection and node supplementation steps only store the response results from GPT-3.5-turbo, with a space complexity of O($1$). Therefore, the overall space complexity of ComGPT is O($nd+kd$).
III Experiments
In this section, we first introduce the experimental settings, including datasets, baselines, implementation details, and evaluation metrics. Then, we give the experimental results and corresponding analyses.
III-A Experimental settings
III-A1 Datasets
TABLE II: Statistics of the datasets.
Networks Nodes Edges $\bar{d}$ $|C|$ $\mu$ $\bar{Q}$
Dolphins 62 159 5.13 31 0.0383 0.3734
Football 115 613 10.66 9.58 0.3638 0.5539
Polbooks 105 441 8.4 35 0.1812 0.4149
Amazon 334863 925872 5.53 30.23 0.1216 0.4441
DBLP 317080 1049866 6.62 53.41 0.2147 0.4453
Five real-world datasets, including Dolphins [25], Football [26], Polbooks [27], Amazon [28], and DBLP [28] are utilized to test the performance of algorithms. Table II shows the detailed information of five real networks, including the number of nodes, the number of edges, the average degree $\bar{d}$, the average community size $|C|$, the mixing parameter $\mu$ [29], and the modularity $Q$ [30].